# Localized RETE for Incremental Graph Queries with Nested Graph Conditions

### v1

@@ Appendix A Technical Details
By Lemma 37, $\sum_{n^{\Phi}\in V^{N^{\Phi}_{Q}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\sum_{n\in V^{N_{Q}}\setminus\{p_{Q}\}}\sum_{m\in\mathcal{C}(n)}|m|+5\cdot\sum_{m\in\mathcal{C}|_{\{p_{Q}\}}}|m|$, since $localize(Q)=localize(N_{Q},p_{Q})$.
In the worst case, the extension points of $(N^{\Phi},p^{\Phi})$ already contain all possible single-vertex matches into $H$ even without the additional input via $RPS_{l}$. Therefore, by the induction hypothesis, it must hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}_{(Q^{\prime},\psi^{\prime})}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{V^{N^{\prime}}}|_{e}$.
From Lemma 12 in [BG24b], it follows that $\sum_{n^{\Phi}_{Q}\in V^{RPS_{l}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q})}|m|\leq 2\cdot|\mathcal{C}|_{\{p_{Q}\}}|_{e}$. Finally, from Theorem 8 and Lemma 12 in [BG24b], it follows that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(p^{\Phi}_{(Q,\psi)}}|m|\leq|\mathcal{C}|_{\{[\ltimes]\}}|_{e}$.
@@
For a nested condition of the form $\psi=\neg\psi^{\prime}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, a request projection structure $RPS_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi^{\prime})})$, and a marking-sensitive anti-join node $p^{\Phi}=[\rhd]^{\Phi}$.
@@
For a nested condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, a request projection structure $RPS_{1}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi_{1})})$, a marking-sensitive semi-join node $[\ltimes]^{\Phi}_{1}$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, a request projection structure $RPS_{2}=RPS([\ltimes]^{\Phi}_{1},N^{\Phi}_{(Q,\psi_{2})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}_{2}$.
@@
From Lemma 12 in [BG24b], it follows that $\sum_{n^{\Phi}_{Q}\in V^{RPS_{1}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q})}|m|\leq 2\cdot|\mathcal{C}|_{\{p_{Q}\}}|_{e}$.
We furthermore know by Theorem 8 that $\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{1})$ can only contain matches for $Q$ that satisfy $\phi_{1}$ and $\mathcal{C}([\ltimes]_{1})$ contains all such matches. From Lemma 12 in [BG24b], it thus follows that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{1})}|m|\leq|\mathcal{C}|_{\{[\ltimes]_{1}\}}|_{e}$. Consequently, it must hold that $\sum_{n^{\Phi}_{Q}\in V^{RPS_{2}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q})}|m|\leq 2\cdot|\mathcal{C}|_{\{[\ltimes]_{1}\}}|_{e}$. By Theorem 8 and Lemma 12 in [BG24b], it also follows that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{2})}|m|\leq|\mathcal{C}|_{\{[\ltimes]_{2}\}}|_{e}$.
@@
Theorem 27.
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and and $\mathcal{C}^{\Phi}_{0}$ the empty configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$. Executing $(N^{\Phi},p^{\Phi})$ via $execute(order^{\Psi}(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(T\cdot|Q_{max}|)$ steps, with $T=\sum_{n\in V^{N}}|\mathcal{C}(n)|$ and $|Q_{max}|$ the size of the largest graph among $Q$ and graphs in $\psi$.
@@
All remaining nodes in $(N^{\Phi},p^{\Phi})$ are then marking-sensitive semi-joins, anti-joins or projections, marking assignments, or marking filters that are only executed once. Moreover, each node in $(N^{\Phi},p^{\Phi})$ can be the dependency of at most one of these semi-joins or anti-joins and at most one of these marking filters. By Lemmata 16, 18, and 19 in [BG24b] and Lemmata 41 and 42, the execution of these nodes is thus in $O((T^{\Phi}_{\Psi}+T^{\Phi}_{RPS}+T^{\Phi}_{Q})\cdot Q_{max})$, with $T^{\Phi}_{\Psi}=\sum_{n\in N^{\Phi}_{\Psi}}|\mathcal{C}(n)|$, where $N^{\Phi}_{\Psi}$ denotes the set of all semi-join and anti-join nodes inserted by $localize^{\Psi}$.
@@
Case 1: Let $m_{a}:Q^{\prime}\rightarrow H$ such that $f\circ m_{K}=m_{a}\circ a\wedge m_{a}\models\psi^{\prime}$. It must then either hold that $m_{a}^{\prime}=iso_{s}\circ m_{a}$ is (1.1) not a valid graph morphism from $Q^{\prime}$ into $H^{\prime}$ or (1.2) $m_{a}^{\prime}\not\models\psi^{\prime}$. In case 1.1, since $iso_{s}$ is an isomorphism, it then follows that $m_{a}(Q^{\prime})\cap H_{p}\neq\emptyset$ and thus $f\circ m_{K}$ is subgraph satisfaction dependent. In case 1.2, it follows that there must be some graph morphism $m_{K_{a}}$ such that $m_{a}=f\circ m_{K_{a}}$ and $m_{a}^{\prime}=g\circ m_{K_{a}}$. Consequently, $m_{a}$ must be subgraph satisfaction dependent for $\psi^{\prime}$ or $m_{a}^{\prime}$ must be subgraph satisfaction dependent for $\psi^{\prime}$ by the induction hypothesis. Since in this case, $g\circ m_{K}=m_{a}^{\prime}\circ a$, it follows that $f\circ m_{K}\text{ is subgraph satisfaction dependent }\vee g\circ m_{K}\text{is subgraph satisfaction dependent}$.
@@
Theorem 29.
Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ an extended graph query. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. It then holds that $\forall m\in\mathcal{M}^{Q}_{H}:m\text{ is subgraph satisfaction dependent }\Rightarrow m\in\mathcal{R}^{\Phi}_{X}(p^{sat},\mathcal{C}^{\Phi})$.
@@
Theorem 30.
Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. Executing $(N^{sat},p^{sat})$ via $O=order^{sat}(N^{sat})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
By the definition of $localize^{\Delta}$, $N^{\Delta}$ consists of $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(N^{Q},p^{Q})$ for a regular RETE net $(N^{Q},p^{Q})$ with height $h$ for the plain graph query $Q$, $[\phi>h]^{\Phi}$, $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$, and $[\cup]^{\Phi}$. In addition, $N^{\Delta}$ comprises $(N^{\Phi}_{(Q,\psi)},p^{\Phi}_{(Q,\psi)})=localize^{\Psi}(Q,\psi)$, $RPS_{l}=RPS([\cup]^{\Phi},N^{\Phi}_{(Q,\psi)})$, $[\ltimes]^{\Phi}$, and $[\xrightarrow{f\circ g^{-1}}]^{\Phi}$.
Let $m\in\mathcal{M}^{Q}_{H}$ with $m\models\psi\wedge(m(Q)\cap H_{p}\neq\emptyset\vee m\text{ is subgraph satisfaction dependent }\vee\exists m^{\prime}\in\mathcal{M}^{Q}_{H^{\prime}}:f\circ g^{-1}\circ m^{\prime}=m\wedge m^{\prime}\text{ is subgraph satisfaction dependent})$. We can then distinguish three cases:
@@
Theorem 32.
Let $(Q,\psi)$ be an extended graph query and $H$ a host graph with relevant subgraph $H_{p}$ and and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$. Executing $(N^{\Delta},p^{\Delta})$ via $O=order^{\Delta}(N^{\Delta})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Delta},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
By the definition of $localize^{\Delta}$, $N^{\Delta}$ consists of $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(N^{Q},p^{Q})$ for a regular RETE net $(N^{Q},p^{Q})$ with height $h$ for the plain graph query $Q$, $[\phi>h]^{\Phi}$, $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$, and $[\cup]^{\Phi}$. In addition, $N^{\Delta}$ comprises $(N^{\Phi}_{(Q,\psi)},p^{\Phi}_{(Q,\psi)})=localize^{\Psi}(Q,\psi)$, $RPS_{l}=RPS([\cup]^{\Phi},N^{\Phi}_{(Q,\psi)})$, $[\ltimes]^{\Phi}$, and $[\xrightarrow{f\circ g^{-1}}]^{\Phi}$.
Analogously, $N^{\Delta^{\prime}}$ consists of $(N^{\Phi^{\prime}}_{Q},p^{\Phi^{\prime}}_{Q})=localize(N^{Q},p^{Q})$, $[\phi>h]^{\Phi^{\prime}}$, $(N^{sat^{\prime}},p^{sat^{\prime}})=localize^{sat}(Q,\psi)$, and $[\cup]^{\Phi^{\prime}}$. In addition, $N^{\Delta^{\prime}}$ comprises $(N^{\Phi^{\prime}}_{(Q,\psi)},p^{\Phi^{\prime}}_{(Q,\psi)})=localize^{\Psi}(Q,\psi)$, $RPS^{\prime}_{l}=RPS([\cup]^{\Phi^{\prime}},N^{\Phi^{\prime}}_{(Q,\psi)})$, $[\ltimes]^{\Phi^{\prime}}$, and $[\xrightarrow{g\circ f^{-1}}]^{\Phi^{\prime}}$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ for host graph $H$ and relevant subgraph $H_{p}$ corresponding to $(N,p)$. It then holds that $\sum_{n^{sat}\in V^{N^{sat}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{sat})}|m|\leq 18\cdot|\mathcal{C}|_{e}$.

### v2

@@ Appendix A Technical Details
Lemma 25 entails that $\sum_{n^{\Phi}\in V^{N^{\Phi}_{Q}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\sum_{n\in V^{N_{Q}}\setminus\{p_{Q}\}}\sum_{m\in\mathcal{C}(n)}|m|+5\cdot\sum_{m\in\mathcal{C}|_{\{p_{Q}\}}}|m|$, since $localize(Q)=localize(N_{Q},p_{Q})$.
In the worst case, the extension points of $(N^{\Phi},p^{\Phi})$ already contain all possible single-vertex matches into $H$ even without the additional input via $RPS^{\infty}_{l}$. Therefore, by the induction hypothesis, it must hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}_{(Q^{\prime},\psi^{\prime})}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{V^{N^{\prime}}}|_{e}$.
From Lemma 12 in [BG24b], it follows that $\sum_{n^{\Phi}_{Q}\in V^{RPS^{\infty}_{l}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q})}|m|\leq 2\cdot|\mathcal{C}|_{\{p_{Q}\}}|_{e}$. Finally, from Theorem 8 and Lemma 12 in [BG24b], it follows that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(p^{\Phi}_{(Q,\psi)}}|m|\leq|\mathcal{C}|_{\{[\ltimes]\}}|_{e}$.
@@
For a nested condition of the form $\psi=\neg\psi^{\prime}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, a request projection structure $RPS^{\infty}_{l}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi^{\prime})})$, and marking-sensitive anti-join node $p^{\Phi}=[\rhd]^{\Phi}$.
@@
For a nested condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, a request projection structure $RPS^{\infty}_{1}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi_{1})})$, a marking-sensitive semi-join node $[\ltimes]^{\Phi}_{1}$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, a request projection structure $RPS^{\infty}_{2}=RPS^{\infty}([\ltimes]^{\Phi}_{1},N^{\Phi}_{(Q,\psi_{2})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}_{2}$.
@@
From Lemma 12 in [BG24b], it follows that $\sum_{n^{\Phi}_{Q}\in V^{RPS^{\infty}_{1}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q})}|m|\leq 2\cdot|\mathcal{C}|_{\{p_{Q}\}}|_{e}$.
We furthermore know by Theorem 8 that $\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{1})$ can only contain matches for $Q$ that satisfy $\phi_{1}$ and $\mathcal{C}([\ltimes]_{1})$ contains all such matches. From Lemma 12 in [BG24b], it thus follows that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{1})}|m|\leq|\mathcal{C}|_{\{[\ltimes]_{1}\}}|_{e}$. Consequently, it must hold that $\sum_{n^{\Phi}_{Q}\in V^{RPS^{\infty}_{2}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q})}|m|\leq 2\cdot|\mathcal{C}|_{\{[\ltimes]_{1}\}}|_{e}$. By Theorem 8 and Lemma 12 in [BG24b], it also follows that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{2})}|m|\leq|\mathcal{C}|_{\{[\ltimes]_{2}\}}|_{e}$.
@@
Theorem 4.8.
(Execution time overhead introduced by RETE net localization via $localize^{\Psi}$ depends on query graph size compared to average match size.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}_{0}$ the empty configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$. Executing $(N^{\Phi},p^{\Phi})$ via $execute(order^{\Psi}(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(T\cdot|Q_{max}|)$ steps, with $T=\sum_{n\in V^{N}}|\mathcal{C}(n)|$ and $|Q_{max}|$ the size of the largest graph among $Q$ and graphs in $\psi$.
@@
All remaining nodes in $(N^{\Phi},p^{\Phi})$ are then marking-sensitive semi-joins, anti-joins or projections, marking assignments, or marking filters that are only executed once. Moreover, each node in $(N^{\Phi},p^{\Phi})$ can be the dependency of at most one of these semi-joins or anti-joins and at most one of these marking filters. By Lemmata 16, 18, and 19 in [BG24b] and Lemmata 29 and 30, the execution of these nodes is thus in $O((T^{\Phi}_{\Psi}+T^{\Phi}_{RPS}+T^{\Phi}_{Q})\cdot Q_{max})$, with $T^{\Phi}_{\Psi}=\sum_{n\in N^{\Phi}_{\Psi}}|\mathcal{C}(n)|$, where $N^{\Phi}_{\Psi}$ denotes the set of all semi-join and anti-join nodes inserted by $localize^{\Psi}$.
@@
Case 1: Let $m_{a}:Q^{\prime}\rightarrow H$ such that $f\circ m_{K}=m_{a}\circ a\wedge m_{a}\models\psi^{\prime}$. It must then either hold that $m_{a}^{\prime}=iso_{s}\circ m_{a}$ is (1.1) not a valid graph morphism from $Q^{\prime}$ into $H^{\prime}$ or (1.2) $m_{a}^{\prime}\not\models\psi^{\prime}$. In case 1.1, since $iso_{s}$ is an isomorphism, it then follows that $m_{a}(Q^{\prime})\cap H_{p}\neq\emptyset$ and thus $f\circ m_{K}$ is subgraph satisfaction dependent. In case 1.2, it follows that there must be some graph morphism $m_{K_{a}}$ such that $m_{a}=f\circ m_{K_{a}}$ and $m_{a}^{\prime}=g\circ m_{K_{a}}$. Consequently, $m_{a}$ must be subgraph satisfaction dependent for $\psi^{\prime}$ or $m_{a}^{\prime}$ must be subgraph satisfaction dependent for $\psi^{\prime}$ by the induction hypothesis. Since in this case, $g\circ m_{K}=m_{a}^{\prime}\circ a$, it follows that $f\circ m_{K}\text{ is subgraph satisfaction dependent }\text{ or }g\circ m_{K}\text{ is subgraph satisfaction dependent}$.
@@
In the case where $a$ is a partial graph morphism from some subgraph $Q_{p}\subseteq Q$ into $Q^{\prime}$, the argumentation works analogously.
@@
Theorem 4.12.
(Consistent configurations for RETE nets localized via $localize^{sat}$ yield all subgraph satisfaction dependent matches.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ an extended graph query, and $\mathcal{C}^{\Phi}$ a consistent configuration for the RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. It then holds that $\forall m\in\mathcal{M}^{Q}_{H}:m\text{ is subgraph satisfaction dependent }\Rightarrow m\in\mathcal{R}^{\Phi}_{X}(p^{sat},\mathcal{C}^{\Phi})$.
@@
Theorem 4.13.
(Execution of localized RETE nets via $order^{sat}$ yields consistent configurations.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. Executing $(N^{sat},p^{sat})$ via $O=order^{sat}(N^{sat})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
By the definition of $localize^{\Delta}$, $N^{\Delta}$ consists of $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(N^{Q},p^{Q})$ for a regular RETE net $(N^{Q},p^{Q})$ with height $h$ for the plain graph query $Q$, $[\phi>h]^{\Phi}$, $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$, and $[\cup]^{\Phi}$. In addition, $N^{\Delta}$ comprises $(N^{\Phi}_{(Q,\psi)},p^{\Phi}_{(Q,\psi)})=localize^{\Psi}(Q,\psi)$, $RPS^{\infty}_{l}=RPS^{\infty}([\cup]^{\Phi},N^{\Phi}_{(Q,\psi)})$, $[\ltimes]^{\Phi}$, and $[\xrightarrow{f\circ g^{-1}}]^{\Phi}$.
Let $m\in\mathcal{M}^{Q}_{H}$ with mψ(m(Q)Hp or m is subgraph satisfaction dependent or
mHQ:fg1m=mm is subgraph satisfaction dependent)
[…]
. We can then distinguish three cases:
@@
Theorem 4.16.
(Execution of localized RETE nets via $order^{\Delta}$ yields consistent configurations.) Let $(Q,\psi)$ be an extended graph query and $H$ a host graph with relevant subgraph $H_{p}$ and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$. Executing $(N^{\Delta},p^{\Delta})$ via $O=order^{\Delta}(N^{\Delta})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Delta},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
By the definition of $localize^{\Delta}$, $N^{\Delta}$ consists of $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(N^{Q},p^{Q})$ for a regular RETE net $(N^{Q},p^{Q})$ with height $h$ for the plain graph query $Q$, $[\phi>h]^{\Phi}$, $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$, and $[\cup]^{\Phi}$, which has a dependency on $p^{sat^{\prime}}_{(Q,\psi)}$ from $(N^{sat^{\prime}}_{(Q,\psi)},p^{sat^{\prime}}_{(Q,\psi)})$. In addition, $N^{\Delta}$ comprises $(N^{\Phi}_{(Q,\psi)},p^{\Phi}_{(Q,\psi)})=localize^{\Psi}(Q,\psi)$, $RPS^{\infty}_{l}=RPS^{\infty}([\cup]^{\Phi},N^{\Phi}_{(Q,\psi)})$, $[\ltimes]^{\Phi}$, and $[\xrightarrow{f\circ g^{-1}}]^{\Phi}$.
Analogously, $N^{\Delta^{\prime}}$ consists of $(N^{\Phi^{\prime}}_{Q},p^{\Phi^{\prime}}_{Q})=localize(N^{Q},p^{Q})$, $[\phi>h]^{\Phi^{\prime}}$, $(N^{sat^{\prime}},p^{sat^{\prime}})=localize^{sat}(Q,\psi)$, and $[\cup]^{\Phi^{\prime}}$ which has a dependency on $p^{sat}_{(Q,\psi)}$ from $(N^{sat}_{(Q,\psi)},p^{sat}_{(Q,\psi)})$. In addition, $N^{\Delta^{\prime}}$ comprises $(N^{\Phi^{\prime}}_{(Q,\psi)},p^{\Phi^{\prime}}_{(Q,\psi)})=localize^{\Psi}(Q,\psi)$, $RPS^{\infty^{\prime}}_{l}=RPS^{\infty}([\cup]^{\Phi^{\prime}},N^{\Phi^{\prime}}_{(Q,\psi)})$, $[\ltimes]^{\Phi^{\prime}}$, and $[\xrightarrow{g\circ f^{-1}}]^{\Phi^{\prime}}$.
@@
(RETE net localization via $localize^{sat}$ introduces only a constant factor overhead on effective configuration size.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ for host graph $H$ and relevant subgraph $H_{p}$ corresponding to $(N,p)$. It then holds that $\sum_{n^{sat}\in V^{N^{sat}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{sat})}|m|\leq 18\cdot|\mathcal{C}|_{e}$.